OpenStudy (anonymous):

If x = 2 + root(3), then find the value of {x(square) - 4x + 1}.

OpenStudy (anonymous):

PLEASE REPLY ASAP. PLEASE. I've got an exam tomorrow.

zepdrix (zepdrix):

Recall that when you expand a second degree binomial you get: `square of the first term` plus `twice the product of both terms` plus `square of the second term`. Example:\[\Large\rm (a+b)^2=a^2+2(ab)+b^2\]Let's see if we can apply this to our problem here :)

zepdrix (zepdrix):

\[\Large\rm (\color{orangered}{x})^2-4(\color{orangered}{x})+1\]We want to evaulate this when, \(\Large\rm \color{orangered}{x=2+\sqrt{3}}\)

zepdrix (zepdrix):

\[\Large\rm (\color{orangered}{2+\sqrt{3}})^2-4(\color{orangered}{2+\sqrt{3}})+1\]

zepdrix (zepdrix):

So expanding the square gives us:\[\Large\rm 2^2+2(2\sqrt3)+(\sqrt3)^2-4(\color{orangered}{2+\sqrt{3}})+1\]Understand how I did that? Square of the first, 2 times the product of both in the middle, and square of the second.

zepdrix (zepdrix):

Distribute the -4 to each term in the orange. Then combine like-terms. Shouldn't be too bad :o

OpenStudy (anonymous):

I really didn't understand. Could you be a little more clear. I understand that (x2 - 4x + 1) is an identity (x - 1)2.

zepdrix (zepdrix):

`I understand that (x^2 - 4x + 1) is an identity (x - 1)^2.` What does this mean? :o

zepdrix (zepdrix):

\[\Large\rm x^2-4x+1 \ne (x-1)^2\]\[\Large\rm x^2-2x+1=(x-1)^2\]

OpenStudy (midhun.madhu1987):

 NOW TRY TO SOLVE IT PLEASE.....